# Global parameterizations of $\pi\pi$ scattering with dispersive constraints: beyond the S0 wave

### v1

@@ Acknowledgements.
Appendix A Threshold parameters
Our aim in this work has been to obtain simple and global parameterizations of partial-wave data covering the energy range from threshold to at least 1.8 GeV, simultaneously consistent with the dispersive constraints. We have made them continuous and with a continuous derivative (except as required by thresholds).
Previous dispersive analyses were more focused on the low-energy region, particularly below 1.1 GeV, and, up to 0.9 GeV, we have basically mimicked them. Of course, to cover a much wider energy region with just one parameterization, and being just a mimic of the CFD at low energies, we have sacrificed the precision attained by the CFD. For the $\pi\pi$ threshold parameters, the old CFD is still more accurate, and even more when combined with sum rules, as done in [54].
Nevertheless, we have checked that our threshold parameters are consistent with the values in [54]. Thus, we have gathered in Table 10 the values of threshold parameters for the S, P, D, F, and G waves that result from our constrained Global Fits. We also provide the values used as input, obtained from the CFD parameterization in [54], except for those of the G waves, which are taken from [73]. In addition, we are listing the best values in [54]. These are the most reliable because they are not only obtained from the CFD but from the use of sum rules, and therefore have a much better precision and are parameterization independent.
@@ Appendix A Threshold parameters
Table 11: S0-wave parameters of the constrained Global Fit I, to be used with the parameterization provided in [68].
Appendix B The S0-wave parameterization
Throughout this paper, we have kept the very same global parameterization of the S0 wave provided in [68]. The reasons are that it was already global, there are no novelties in the data for this wave (contrary to the P wave), the existing parameterizations were sufficiently flexible, and the $K\bar{K}$ channel dominates the opening of the inelasticity. However, since we have changed the other waves, when imposing the FDRs we have allowed small variations in the S0-wave parameters within uncertainties. As a result, these parameters change slightly, particularly, those affecting the S0 wave above 1.4 GeV. We have gathered the new parameters in Tables 1112 and 13. Of course, we use them with the very same parameterization provided in [68] (check the erratum too).
@@ Appendix B The S0-wave parameterization
Table 12: S0-wave parameters of the constrained Global Fit II, to be used with the parameterization provided in [68].
@@
Table 13: S0-wave parameters of the constrained Global Fit III, to be used with the parameterization provided in [68].
In Fig. 19 we show our slightly updated Global Fit I S0 wave compared with the Global I in [68]. As expected, the changes are very small and affect mostly the region above 1.4 GeV. A similar situation occurs with Global Fits II and III and their counterparts in [68]
Finally, we show the three Global Fits in Fig. 20. As usual, they are fairly compatible up to 1.4 GeV, but differ widely above. For us here, it is interesting to note that between 1 and 1.10 GeV, the data Solutions II and III of Hyams et al. 75 [26] prefer an elasticity $\eta\simeq 0.75$. This value corresponds to the “non-dip” solution, which is strongly disfavored by Roy and GKPY analyses [54, 56]. Thus, in this work, as in [68], we are imposing in the Global Fits the “dip-solution” from the CFD in [54]. As a result, the Global Fits II and III have such a dip, although their data do not call for it, while still describing the rest of data from [26] above 1.15 GeV.
Figure 19: S0-wave phase shift (top) and elasticity (bottom). We show the result of our dispersively constrained Global Fit I versus the “Old Global I” parameterization from [68]. Above 1.6 GeV (shaded region) there are no dispersive constraints. The data comes from the Solution B of Grayer et al. [25], Kaminski et al. [28], and NA48/2 [31].

### v2

@@ Appendix A Threshold parameters
Appendix B The S0-wave parametrization
Throughout this paper, we have kept the very same global parametrization of the S0 wave provided in [68]. The reasons are that it was already global, there are no novelties in the data for this wave (contrary to the P wave), the existing parametrizations were sufficiently flexible, and the $K\bar{K}$ channel dominates the opening of the inelasticity. However, since we have changed the other waves, when imposing the FDRs we have allowed small variations in the S0-wave parameters within uncertainties. As a result, these parameters change slightly, particularly, those affecting the S0 wave above 1.4 GeV. We have gathered the new parameters in Tables 1112 and 13. Of course, we use them with the very same parametrization provided in [68] (check the erratum too).
@@ Appendix B The S0-wave parametrization
Table 11: S0-wave parameters of the constrained Global Fit I, to be used with the parametrization provided in [68].
In Fig. 19, we show our slightly updated Global Fit I S0 wave compared with the Global I in [68]. As expected, the changes are very small and affect mostly the region above 1.4 GeV. A similar situation occurs with Global Fits II and III and their counterparts in [68].
Finally, we show the three Global Fits in Fig. 20. As usual, they are fairly compatible up to 1.4 GeV but differ widely above. For us here, it is interesting to note that between 1 and 1.10 GeV, the data Solutions II and III of Hyams et al. 75 [26] prefer an elasticity $\eta\simeq 0.75$. This value corresponds to the “non-dip” solution, which is strongly disfavored by Roy and GKPY analyses [54, 56]. Thus, in this work, as in [68], we are imposing in the Global Fits the “dip-solution” from the CFD in [54]. As a result, the Global Fits II and III have such a dip, although their data do not call for it, while still describing the rest of data from [26] above 1.15 GeV.
@@
Table 12: S0-wave parameters of the constrained Global Fit II, to be used with the parametrization provided in [68].
@@
Table 13: S0-wave parameters of the constrained Global Fit III, to be used with the parametrization provided in [68].
Figure 19: S0-wave phase shift (top) and elasticity (bottom). We show the result of our dispersively constrained Global Fit I versus the “Old Global I” parametrization from [68]. Above 1.6 GeV (shaded region) there are no dispersive constraints. The data comes from the Solution B of Grayer et al. [25], Kaminski et al. [28], and NA48/2 [31].